OpenStudy (rainbow_rocks03):

Choose the missing number to make equivalent fractions. 5/6 = 20/___ ? A. 21 B. 22 C. 24 D. 30

OpenStudy (rainbow_rocks03):

PLZ HELP and tell me how.

OpenStudy (haseeb96):

put one by one all the option and then solve them and see which option answer is matched with it .

geerky42 (geerky42):

Well, when you do something to numerator, like multiplying by something, you must do same in denominator, right?

OpenStudy (rainbow_rocks03):

so 5*6 & 20*6 ?

OpenStudy (haseeb96):

No first we will try to put 22

OpenStudy (rainbow_rocks03):

or 20*6=___ divided by 6=___

OpenStudy (haseeb96):

put its answer dont match with the given answer now put 24 when we will solve this 20/24 you will get 5/6 So C is the correct answer

geerky42 (geerky42):

No, for example in \(\dfrac{1}{2}\), when you multiply numerator by 2, you must do same for denominator, so value of fraction won't change. \(\dfrac{1\times2}{2\times2}\)

geerky42 (geerky42):

Here, you have \(\dfrac{5}{6}\) What should you do to numerator to get to 20?

OpenStudy (rainbow_rocks03):

Is that really the answer because I need to get an 80% or higher?

OpenStudy (rainbow_rocks03):

never mind when I need more help plz help

geerky42 (geerky42):

Ok you only want answer. Sorry then.